# Anton Alexa

### v3

@@ Abstract
We establish a boundary delocalization principle for high–frequency Dirichlet eigenfunctions on smooth strictly convex domains. The main result is an exclusion of persistent boundary concentration at the level of individual eigenmodes relative to short spectral packets of sublinear length. Quantitatively, we compare boundary energies of single eigenfunctions to packet sums over frequency windows of size $N_{k}=o(k)$, without asserting any asymptotic gain in magnitude. The main mode–to–packet estimate relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain a packet zero–mean cancellation estimate. This mode–to–packet comparison is independent of eigenvalue monotonicity and is stable under eigenvalue crossings.
@@ 1 Introduction
Related results on boundary delocalization and restriction estimates typically rely on microlocal analysis, defect measures, or dynamical properties of the billiard flow [5, 6, 12, 7, 1]. For general background on boundary eigenfunctions and related phenomena [14, 15]. In contrast, the main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain packet zero–mean cancellation.
This paper shows that, on strictly convex domains, persistent boundary concentration at the level of a single mode is excluded when one compares it to a spectral packet of sublinear length. We establish a packet–level delocalization principle: for any sublinear window $N_{k}=o(k)$, the ratio between a single boundary energy and the packet sum tends to zero. The proof uses the Rellich identity. In addition, we rule out collective boundary concentration: no positive fraction of modes inside a sublinear spectral packet can exhibit a persistent one–sided boundary bias with a fixed geometric weight of zero mean, in the spirit of averaged spectral quantities; [11, 10]. For the multi–mode statement we assume a slightly stronger growth condition on $N_{k}$ when $d\geq 3$.
@@
In particular, without averaging over a growing spectral window, a single eigenmode may contribute a non–negligible fraction of the packet boundary energy. Thus the requirement $N_{k}\to\infty$ represents the minimal averaging scale necessary to enforce mode–to–packet delocalization.
@@ 2 Preliminaries and notation
All arguments are formulated at the level of boundary energy densities $\rho_{k}=|\partial_{n}u_{k}|^{2}$ and spectral sums. No ordering or differentiability of individual eigenvalues is required, and possible eigenvalue crossings do not affect the packet–level estimates.
@@ 4 Packet scales and a contextual Weyl remark
Remark 4.5 (Packet size condition).
For $d=2$ the condition $N_{k}/k^{1-2/d}\to\infty$ reduces to $N_{k}\to\infty$. For $d\geq 3$ it requires only mildly larger sublinear packets.
@@
All constants appearing in Lemma 3.1 and Proposition 4.7 are for the fixed strictly convex domain $\Omega$.
@@ Conclusion
The main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law; integrated boundary Weyl asymptotics for the Dirichlet Laplacian [2, 4, 8] are recorded for context. In particular, no microlocal defect measures, dynamical assumptions on the billiard flow, or ergodicity hypotheses are required. The argument is robust under eigenvalue crossings and depends only on global spectral information.
From a conceptual viewpoint, the result provides a mode–to–packet comparison mechanism on the boundary, analogous to bulk delocalization phenomena but valid in a purely deterministic and geometric setting. The sublinear packet scale reflects the minimal averaging required to control single–mode contributions using Rellich-type bounds. In addition, we obtain a multi–mode delocalization principle: no positive fraction of modes in a sublinear packet can maintain strong boundary correlation with a fixed zero–mean weight.

### v4

@@ Abstract
We establish a boundary delocalization principle for high–frequency Dirichlet eigenfunctions on smooth strictly convex domains. The main result is an exclusion of persistent boundary concentration at the level of individual eigenmodes relative to short spectral packets of sublinear length. Quantitatively, we compare boundary energies of single eigenfunctions to packet sums over frequency windows of size $N_{k}=o(k)$, without asserting any asymptotic gain in magnitude. The main mode–to–packet estimate relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain a packet zero–mean cancellation estimate at the scale $N_{k}/k^{1-2/d}\to\infty$. This mode–to–packet comparison is independent of eigenvalue monotonicity and is stable under eigenvalue crossings.
@@ 1 Introduction
Related results on boundary delocalization and restriction estimates typically rely on microlocal analysis, defect measures, or dynamical properties of the billiard flow [5, 6, 12, 7, 1]. For general background on boundary eigenfunctions and related phenomena [14, 15]. Classical boundary quantum ergodicity and boundary-trace completeness results include [16, 17, 18, 19]. For recent broader developments in related semiclassical restriction and eigenfunction analysis, [20, 21]. In contrast, the main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain packet zero–mean cancellation. Our approach is complementary in spirit to microlocal and dynamical studies of boundary eigenfunctions [5, 6, 12, 7, 1, 14, 15]. We establish a deterministic mode–to–packet comparison principle based on global spectral identities: Rellich bounds for single modes and boundary local Weyl cancellation in zero–mean pairings for the multi–mode statement [2, 4, 8]. The primary motivation for this packet–level viewpoint is structural stability under spectral deformations. In settings with shape perturbations or geometric flows, eigenvalue crossings and changing multiplicities are typical. Since the estimates below are formulated via spectral sums, they are invariant under changes of orthonormal bases inside eigenspaces and remain stable across such crossings.
This paper shows that, on strictly convex domains, persistent boundary concentration at the level of a single mode is excluded when one compares it to a spectral packet of sublinear length. We establish a packet–level delocalization principle: for any sublinear window $N_{k}=o(k)$, the ratio between a single boundary energy and the packet sum tends to zero. The proof uses the Rellich identity. In addition, we rule out collective boundary concentration: no positive fraction of modes inside a sublinear spectral packet can exhibit a persistent one–sided boundary bias with a fixed geometric weight of zero mean, in the spirit of averaged spectral quantities; [11, 10]. For the multi–mode statement we assume
$\frac{N_{k}}{k^{1-2/d}}\to\infty,$ (1)
which is the methodological threshold for packet cancellation in the present Rellich + boundary Weyl framework.
@@
In particular, without averaging over a growing spectral window, a single eigenmode may contribute a non–negligible fraction of the packet boundary energy. Thus the requirement $N_{k}\to\infty$ represents the minimal averaging scale necessary to enforce single–mode to packet delocalization.
@@ 2 Preliminaries and notation
All arguments are formulated at the level of boundary energy densities $\rho_{k}=|\partial_{n}u_{k}|^{2}$ and spectral sums. No ordering or differentiability of individual eigenvalues is required, possible eigenvalue crossings do not affect the packet–level estimates, and the bounds are invariant under orthonormal basis changes in eigenspaces.
@@ 4 Packet scales and a contextual Weyl remark
Proposition 4.5 (Methodological threshold for packet cancellation).
Under the assumptions of Lemma 4.3, one has
$\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\lesssim\frac{k^{1-2/d}}{N_{k}}.$ (27)
In particular, packet-level cancellation follows whenever
$\frac{N_{k}}{k^{1-2/d}}\to\infty.$ (28)
If $N_{k}\lesssim k^{1-2/d}$ along a subsequence, the present method (Rellich identity combined with the boundary Weyl law) does not suffice to ensure cancellation.
Proof 4.6.
The estimate
$\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\lesssim\frac{k^{1-2/d}}{N_{k}}$ (29)
is exactly (26) from the proof of Lemma 4.3. Hence cancellation follows when $N_{k}/k^{1-2/d}\to\infty$. If $N_{k}\lesssim k^{1-2/d}$ along a subsequence, this upper bound is not forced to vanish, so cancellation cannot be concluded by this method alone. ∎
Remark 4.7 (Threshold interpretation).
The exponent 12/d1-2/d comes from boundary Weyl fluctuations after cancellation of the principal term in the zero-mean pairing $\int_{\partial\Omega}Q_{\Lambda}w\,d\sigma$. For $d=2$, the condition $N_{k}/k^{1-2/d}\to\infty$ reduces to $N_{k}\to\infty$; for $d\geq 3$, it requires mildly larger sublinear packets.
@@
All constants appearing in Lemma 3.1 and Propositions 4.5 and 4.9 are for the fixed strictly convex domain $\Omega$.
@@ Conclusion
The main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law; integrated boundary Weyl asymptotics for the Dirichlet Laplacian [2, 4, 8] are recorded for context. The argument is robust under eigenvalue crossings and depends only on global spectral information. This structural invariance makes the packet framework natural for deformation-type questions, where multiplicities and mode labeling may vary.
From a conceptual viewpoint, the result provides a mode–to–packet comparison mechanism on the boundary, analogous to bulk delocalization phenomena but valid in a purely deterministic and geometric setting. For single–mode comparison, the sublinear packet regime with $N_{k}\to\infty$ and $N_{k}=o(k)$ is sufficient. Notably, the single–mode packet delocalization estimate is obtained without using local Weyl asymptotics. For packet zero–mean cancellation and multi–mode bias exclusion, the present method identifies the threshold scale $N_{k}\gg k^{1-2/d}$. The packet–level perspective developed here is particularly suited to averaged spectral functionals, where single–eigenfunction control is often too weak. In addition, we obtain a multi–mode delocalization principle: no positive fraction of modes in a sublinear packet can maintain strong boundary correlation with a fixed zero–mean weight.